SUBJECTS
|
BROWSE
|
CAREER CENTER
|
POPULAR
|
JOIN
|
LOGIN
Business Skills
|
Soft Skills
|
Basic Literacy
|
Certifications
About
|
Help
|
Privacy
|
Terms
|
Email
Search
Test your basic knowledge |
GMAT Quantitative General
Start Test
Study First
Subjects
:
gmat
,
math
Instructions:
Answer 50 questions in 15 minutes.
If you are not ready to take this test, you can
study here
.
Match each statement with the correct term.
Don't refresh. All questions and answers are randomly picked and ordered every time you load a test.
This is a study tool. The 3 wrong answers for each question are randomly chosen from answers to other questions. So, you might find at times the answers obvious, but you will see it re-enforces your understanding as you take the test each time.
1. How to find the slope.
0.15n + 0.08(5) = 0.1(n+5)
The probability of event occurring is...
y2 - y1 / x2 - x1
$11 - 025
2. Number added or deleted
(total distance) / (total time)
If the outcome of one event affects the outcome of the other event.
Total = mean x (number of terms) Number deleted = (original total) - (new total) Number added = (new total) - (original total)
gcd(m,n)*lcm(m,n) = mn
3. The average of consecutive numbers
3 - 6 - 9 - 12
The average of a set of evenly spaced consecutive numbers is the average of the smallest and largest numbers in the set.
Immediately UNFACTOR or vice versa
347
4. Probability and Geometry.
Sum of digits is multiple of 3
$11 - 025
x(sq. rt 3) - x - 2x
If a point is chosen at random within a space with an area - volume - or length of Y and a space with a respective area - volume - or length of X lies within Y - the probability of choosing a random point within Y is the area - volume - or length of
5. In general - difficult questions require how many steps to solve?
at least 3 steps
Number is a multiple of 3 and 2
1/16
Minor arc = 2(inscribed angle)
6. 2n+1 - 2n+3 - 2n+5
Odd
1.7
P(E) + P(F) - P(E and F)
1. Start by writing each number as a product of primes. 2. Write so that each new prime factor begins in the same place. 3. Lowest common multiple is found by multiplying all factors in either list.
7. 45-45-90 triangle basic lengths of sides
Last two digits are multiple of 4 or the number can be divided by 2 twice.
x - x - x(sq. rt 2)
Sum of digits is multiple of 3 - last two digits multiple of 4.
Principal (1 + interest/number times compounded)^(t)(n)
8. Sq. rt(2)
180(n-2)
1.4
Find simple interest then look for the answer that is a little bigger
12.5%
9. 2nd Rule of Probability: P(E) = 1 - P(not E)
Even
(sum of bases)(height) / 2
The probability of an event occurring plus the probability of the event not occurring = 1
The total amount before any deductions
10. 1. A and B < A or B 2. A or B > Individual probabilities of A - B 3. P(A and B) = P(A) x P(B) <-- 'fewer options' 4. P(A or B) = P(A) + P(B) <-- 'more options' - Probability of multiple events rules.
1. Start by writing each number as product of primes. 2. Write so that each new prime factor begins in the same place. 3. Greatest Common Factor is found by multiplying all factors appearing in BOTH lists
1st Rule of Probability: Basic Rule is what?
The total amount before any deductions
0.15n + 0.08(5) = 0.1(n+5)
11. Compound interest rule
Find simple interest then look for the answer that is a little bigger
P(E)P(F)
always try to factor
83.3%
12. Simple Interest Formula (remember this is the total amount of money in the bank after the interest is earned)
A = P(1 + r) ^n
market value
1. Start by writing each number as product of primes. 2. Write so that each new prime factor begins in the same place. 3. Greatest Common Factor is found by multiplying all factors appearing in BOTH lists
Consider work done in one hour. Inverse of the time it takes everyone working together = Sum of the inverse of the times it would take each person working individually.
13. Multiplication principle
Exterior angle d is equal to the sum of the two remote interior angles a and b
the probability of event A AND event B occurring is the probability of event A times the probability of event B - given that A has already occurred.
if a first object may be chosen in m ways and a second object may be chosen in n ways - then there are mn ways of choosing both objects
1/16
14. How to check whether a number is a multiple of 4.
The number of ways independent events can occur together can be determined by multiplying together the number of possible outcomes for each event.
1
Last two digits are multiple of 4 or the number can be divided by 2 twice.
1. Start by writing each number as product of primes. 2. Write so that each new prime factor begins in the same place. 3. Greatest Common Factor is found by multiplying all factors appearing in BOTH lists
15. Think of averages as what? The average of 3 - 4 - 5 - and x is 5. What is x? 3 is 2 less than 5 4 is 1 less than 5 - 5 is the average - x = 5 + 3 = 8
1 - P(E)
Balancing
Total = mean x (number of terms) Number deleted = (original total) - (new total) Number added = (new total) - (original total)
180(n-2)
16. Volume of a sphere
If a point is chosen at random within a space with an area - volume - or length of Y and a space with a respective area - volume - or length of X lies within Y - the probability of choosing a random point within Y is the area - volume - or length of
16.6%
(total distance) / (total time)
4/3 TT r ^3
17. 3rd Rule of Probability: Conditional Probability
D or E
Principal (1 + interest/number times compounded)^(t)(n)
Consider work done in one hour. Inverse of the time it takes everyone working together = Sum of the inverse of the times it would take each person working individually.
the probability of event A AND event B occurring is the probability of event A times the probability of event B - given that A has already occurred.
18. x^r/s = ?
12.5%
1.7
1.4
s Sq. rt (x^r)
19. In general - medium questions require how many steps to solve?
2 steps
A = P(1 + r) ^n
1 - P(E)
1.4
20. (1/4)^2
1/16
x(sq. rt 3) - x - 2x
$11 - 025
Group 1 + Group 2 + Neither - Both = Total
21. Price purchased for by wholesaler
2 steps
3 - 6 - 9 - 12
12^3
Purchase price
22. Inscribed Angle - Minor Arc
Minor arc = 2(inscribed angle)
Even
if a first object may be chosen in m ways and a second object may be chosen in n ways - then there are mn ways of choosing both objects
(total A) / (total B)
23. Quadratic formula
-b +- sq. rt(b^2 - 4ac) / 2a
(total distance) / (total time)
(amount of change) / (original amount)
Last two digits are multiple of 4 or the number can be divided by 2 twice.
24. Price sold for by retailer (after markup)
market value
1st Rule of Probability: Basic Rule is what?
1. Start by writing each number as product of primes. 2. Write so that each new prime factor begins in the same place. 3. Greatest Common Factor is found by multiplying all factors appearing in BOTH lists
Immediately try factoring/simplifying when possible
25. To determine the number of integers less than 5000 that are evenly divisible by 15...?
If a point is chosen at random within a space with an area - volume - or length of Y and a space with a respective area - volume - or length of X lies within Y - the probability of choosing a random point within Y is the area - volume - or length of
0.15n + 0.08(5) = 0.1(n+5)
if a first object may be chosen in m ways and a second object may be chosen in n ways - then there are mn ways of choosing both objects
Divide 4999 by 15 => 333 integers
26. Net
Any multiplication involving an even number creates an even product.
The probability of an event occurring plus the probability of the event not occurring = 1
The average of a set of evenly spaced consecutive numbers is the average of the smallest and largest numbers in the set.
The amount after deductions
27. Simple probability
(# of favorable outcomes) / (# of possible outcomes)
Gross Profit = Selling Price - Cost
The probability of event occurring is...
Check each prime number up to the approximate square root of the number. If you haven't found a number less than or equal to the square root of the number - then the number is prime.
28. Dependent events: When are two events said to be dependent events?
at least 3 steps
For a fixed distance - the average speed is inversely related to the amount of time required to make the trip.
Figure out the probability for each individual event. Multiply the individual probabilities together.
If the outcome of one event affects the outcome of the other event.
29. Percent Formula
P(event NOT occurring) = 1 - P(event occurring)
Odd
p/100 = is/of
Odd numbers only have ___________
30. 2n - 2n+2 - 2n+4
Even
The probability of event A OR B occurring is the probability of event A occurring plus the probability of event B occurring minus the probability of both events occurring. P(A or B) = P(A) +P(B) - P(A and B)
For a fixed distance - the average speed is inversely related to the amount of time required to make the trip.
principle (interest rate - in decimal form) (time - in years)
31. Since Mieko's average speed was 3/4 of Chan's - her time was 4/3 as long.
For a fixed distance - the average speed is inversely related to the amount of time required to make the trip.
Principal (1 + interest/number times compounded)^(t)(n)
The average of a set of evenly spaced consecutive numbers is the average of the smallest and largest numbers in the set.
1. Start by writing each number as a product of primes. 2. Write so that each new prime factor begins in the same place. 3. Lowest common multiple is found by multiplying all factors in either list.
32. The number of outcomes that result in A divided by the total number of possible outcomes.
The probability of an event occurring plus the probability of the event not occurring = 1
(n-1)!
The probability of event occurring is...
The average of a set of evenly spaced consecutive numbers is the average of the smallest and largest numbers in the set.
33. Odd and Even rule.
Any multiplication involving an even number creates an even product.
Sum of digits is multiple of 3
Find simple interest then look for the answer that is a little bigger
Group 1 + Group 2 + Neither - Both = Total
34. 4th rule of Probability
The probability of event A OR B occurring is the probability of event A occurring plus the probability of event B occurring minus the probability of both events occurring. P(A or B) = P(A) +P(B) - P(A and B)
Organize into a grid.
Immediately UNFACTOR or vice versa
Minor arc = 2(inscribed angle)
35. If $10 -000 is invested at 10% annual interest - compounded semi-annually - what is the balance after 1 year?
P(E)P(F)
347
4/3 TT r ^3
$11 - 025
36. Trial Problems: look at the probability of NOT OCCURRING
P(event NOT occurring) = 1 - P(event occurring)
1st Rule of Probability: Basic Rule is what?
Number is a multiple of 3 and 2
If a point is chosen at random within a space with an area - volume - or length of Y and a space with a respective area - volume - or length of X lies within Y - the probability of choosing a random point within Y is the area - volume - or length of
37. What to do with equations that have fractions
If the outcome of one event affects the outcome of the other event.
1.7
Immediately try factoring/simplifying when possible
at least 3 steps
38. How to check whether a number is a multiple of 12.
180(n-2)
To find the number of distinct permutations of a set of items with indistinguishable ('repeat') items - divide the factorial of the items in the set by the product of the factorials of the number of indistinguishable elements.
Sum of digits is multiple of 3 - last two digits multiple of 4.
Find simple interest then look for the answer that is a little bigger
39. Combined Events: E or F
12^3
P(E) + P(F) - P(E and F)
always try to factor
Find simple interest then look for the answer that is a little bigger
40. If you have to guess in a problem - which ones should you guess? Especially if you have to plug numbers.
______ |m-n|
D or E
Immediately try factoring/simplifying when possible
3 - 6 - 9 - 12
41. The average of 5 numbers is 2. After one number is deleted - the new average is -3. What number was deleted?
(n-1)!
16.6%
12^3
22
42. 1/6 = what %
The probability of an event occurring plus the probability of the event not occurring = 1
16.6%
The probability of event A OR B occurring is the probability of event A occurring plus the probability of event B occurring minus the probability of both events occurring. P(A or B) = P(A) +P(B) - P(A and B)
12.5%
43. How do you multiply roots together.
multiply or divide the numbers outside the radical signs - then the numbers inside the radical signs
Balancing
(# of favorable outcomes) / (# of possible outcomes)
83.3%
44. 3^3 x 4^3 = ?
The probability of an event occurring plus the probability of the event not occurring = 1
12^3
If the outcome of one event affects the outcome of the other event.
(# of favorable outcomes) / (# of possible outcomes)
45. How to find all divisors of a number
Any multiplication involving an even number creates an even product.
Find all prime factors
Total = mean x (number of terms) Number deleted = (original total) - (new total) Number added = (new total) - (original total)
12.5%
46. 0! = ?
1.4
1
The probability of event A OR B occurring is the probability of event A occurring plus the probability of event B occurring minus the probability of both events occurring. P(A or B) = P(A) +P(B) - P(A and B)
D or E
47. How to check whether a number is a multiple of 3.
Sum of digits is multiple of 3
2 steps
P(E)P(F)
(x-n(n)y-n)
48. Combined Events: E and F
P(E)P(F)
Figure out the probability for each individual event. Multiply the individual probabilities together.
83.3%
y2 - y1 / x2 - x1
49. Combined Events: Not E = P(not E) = ?
Find simple interest then look for the answer that is a little bigger
Immediately try factoring/simplifying when possible
Sum of digits is multiple of 3 - last two digits multiple of 4.
1 - P(E)
50. Properties of 0
P(event NOT occurring) = 1 - P(event occurring)
Even integer. Neither positive nor negative. Multiple of every number. Not a factor of any number.
market value
| A union B| = |A| + |B| - |A intersect B|